# Prevention is better than cure? Feedback from high specific energy winds in cosmological simulations with Arkenstone

### v1

@@ Abstract
We deploy the new Arkenstone galactic wind model in cosmological simulations for the first time, allowing us to robustly resolve the evolution and impact of high specific energy winds. In a (25 $h^{-1}$ Mpc)3 box we perform a set of numerical experiments that systematically vary the mass and energy loadings of such winds, finding that their energy content is the key parameter controlling the stellar to dark matter mass ratio. Increasing the mass loading, at fixed energy, actually results in mildly enhanced star formation, counter to prevailing wisdom but in agreement with recent analytic models. Of the simple parametrisations that we test, we find that an energy loading that scales inversely with halo mass best matches a wide range of observations and can do so with mass loadings drastically lower than those in most previous cosmological simulations. In this scenario, much less material is ejected from the interstellar medium. Instead, winds both heat gas in the circumgalactic medium, slowing infall onto the galaxy, and also drive shocks beyond the virial radius, preventing accretion onto the halo in the first place. We have not yet tied the mass and energy loadings to high-resolution simulations (a key goal of the Learning the Universe collaboration); however, we can already report that a much lower fraction of the available supernova energy is needed in preventative galaxy regulation than required by ejective wind feedback models such as IllustrisTNG.
@@
pubyear: 2024pagerange: Prevention is better than cure? Feedback from high specific energy winds in cosmological simulations with ArkenstonePrevention is better than cure? Feedback from high specific energy winds in cosmological simulations with Arkenstone
@@ 1 Introduction
To first order, in the $\Lambda$CDM cosmological model, galaxy formation just involves the gravitational collapse of dark matter (DM) hierarchically into DM haloes, with subsequent condensation of baryons into their gravitational potential followed by the formation of stars (Rees & Ostriker, 1977; Silk, 1977; White & Rees, 1978; White & Frenk, 1991). However, reality is likely much more complicated, due to the complex interplay of accretion onto galaxies and their interstellar medium (ISM), along with outflows of mass and energy driven by feedback processes. This feedback is thought to be essential, as simulations that do not include it significantly overproduce stars, leading to very massive, compact galaxies (e.g. Navarro & White, 1994; Navarro et al., 1995).
@@
Feedback is generally believed to come from two primary sources: black holes (BHs) and stars (see e.g. Heckman & Best, 2023). In lower mass haloes, below approximately $10^{12}$ M, the dominant effect is thought to be stellar feedback, from stellar winds, radiation, and supernovae, while massive galaxies are regulated by BH feedback (see e.g. Somerville & Davé, 2015). At all mass scales, regulation of star formation occurs ultimately by reducing the efficiency of the conversion of gas into stars. On small spatial scales, energy injection in the galaxy causes heating and drives turbulence in the ISM, disrupting star formation (e.g., Ostriker et al., 2010; Hopkins et al., 2011; Faucher-Giguère et al., 2013; Krumholz et al., 2018; Ostriker & Kim, 2022). The key to the long-term regulation of star formation however lies in the larger scale baryon cycle, in which galactic winds play a key role (e.g., Strickland & Stevens, 2000; Oppenheimer & Davé, 2008; Hopkins et al., 2012; Anglés-Alcázar et al., 2017; Wright et al., 2024).
@@
Winds can act as feedback in two ways. They can remove material from the ISM, reducing the amount of gas available to form stars and carrying mass and metals out into the circumgalactic medium (CGM) – ejective feedback. They can also slow or stop gas from accreting onto the galaxy to replenish the fuel for star formation, via the injection of energy into the CGM (or beyond) – preventative feedback. We can characterise winds with two key parameters: the mass loading ($\eta_{\rm M}$, the ratio of the wind mass flux and the star formation rate) and the energy loading ($\eta_{\rm E}$, the ratio of the wind energy flux and the rate at which energy is released during star formation and stellar evolution). We will discuss these parameters extensively in this work.
@@
A wealth of observational evidence exists for outflowing material around many galaxies (see Rupke, 2018, for a review), including from spectroscopic detections of outflows themselves and metal pollution of the CGM at large distances from the central galaxy (e.g. Prochaska et al., 2017; Guo et al., 2020; Zabl et al., 2021). These outflows are ubiquitously found to be multiphase, with molecular, neutral and ionised gas found to co-exist using probes across the electromagnetic spectrum (e.g. Rupke, 2018; Thompson & Heckman, 2024). Despite the widespread detection of outflows around galaxies, estimating the mass and energy loading values is very difficult – especially in the hottest gas phases. Where such data exists, mass fluxes are reliably found to be low relative to the star formation rate (e.g. McQuinn et al., 2019; Concas et al., 2022; Marasco et al., 2022; Kado-Fong et al., 2024) and the energy flux is found to be comparable to the energy injected by supernovae (e.g. Strickland & Heckman, 2009).
@@
Most modern cosmological box simulations tune the parameters of their models to match particular observables (e.g. the stellar mass – halo mass relation). After this process, the mass and energy loading factors for stellar feedback in most cosmological simulations are typically much higher than what is measured in resolved ISM simulations (e.g. Li & Bryan, 2020). Different simulations can have drastically different predictions for secondary properties of galaxies, most notably in the CGM (Fielding et al., 2017; Davies et al., 2019, 2020; Fielding et al., 2020; Kelly et al., 2022; Ayromlou et al., 2023; Wright et al., 2024). Resolution also affects the properties of the simulated CGM, further complicating the matter, though some recent progress has been made using simulations that specifically focus computational resources on the CGM (e.g. Hummels et al., 2019; Peeples et al., 2019; van de Voort et al., 2019; Suresh et al., 2019; Bennett & Sijacki, 2020; Ramesh & Nelson, 2024). How much of the discrepancies between simulations are due to the numerical implementation of subgrid models is unclear; until now it has not been possible to study the regulatory effect of high specific energy, low mass winds on a large population of galaxies.
@@
Simulating high specific energy winds is challenging in cosmological simulations (see section 2 of Smith et al., 2024a, for a detailed description). Hot, fast winds with low mass fluxes have lower densities, leading to coarser spatial resolution in a Lagrangian code (or with a quasi-Lagrangian refinement strategy in an Eulerian scheme, as is typical in cosmological volumes). Even with the highest mass resolution of current cosmological box simulations ($\sim\!10^{5}$ M) such a wind can become completely unresolved, with single resolution elements being larger than the galaxy launching them (see fig. 1 of Smith et al., 2024a).
@@
To address these issues, Arkenstone  a new subgrid model for multiphase winds, was developed (Smith et al., 2024a, b). The full model launches winds containing both a high specific energy component (Arkenstone-Hot, Smith et al., 2024a, henceforth S24) and an unresolved cold cloud component (Arkenstone-Cold, Smith et al., 2024b), each with their own mass and energy loadings. With the full model, we can explore a huge area of previously unexplored parameter space to investigate the balance between ejective and preventative feedback in galaxy formation. The ultimate aim is to use parameters calibrated in highly resolved ISM simulations as inputs to the Arkenstone model, to remove the calibration step common to most cosmological simulations (e.g. Kim et al., 2020b) – key aims of the SMAUG [^1] https://www.simonsfoundation.org/flatiron/center-for-computational-astrophysics/galaxy-formation/smaug/ (Simulating Multiscale Astrophysics to Understand Galaxies) and Learning the Universe [^2] http://learning-the-universe.org collaborations.
@@
In this first paper applying the Arkenstone model to cosmological simulations we focus only on the hot component of winds, allowing us to study galaxy formation in a regime where preventative feedback dominates and little mass is ejected from the ISM. We vary the key parameters of these hot winds – mass and energy loadings – to study how the properties of the galaxy population respond. The paper is structured as follows: In Section 2 we describe in detail the Arkenstone model and the simulations in this paper. Sections 3.1 and 3.2 show properties of the galaxies and their gaseous haloes in our simulations at $z=0$. Section 3.4 then looks at a subset of these quantities at $z=2$ to demonstrate how Arkenstone regulates star formation through preventative, rather than ejective feedback. We look at the redshift evolution of star formation in Section 3.5, discuss comparisons to recent analytic, semi-analytic and simulation work in Section 4, and summarise our results in Section 5.
@@ 2.1 Simulation set-up
The simulations in this paper use Arepo (Springel, 2010), a cosmological code that incorporates gravity using a TreePM scheme and solves the equations of hydrodynamics on a quasi-Lagrangian, unstructured, moving Voronoi mesh, where mesh generating points move with the fluid velocity (with small corrections in selected cells to promote a regular mesh geometry). In Arepo, a default refinement scheme keeps gas cells within a factor of two of a given mass resolution by merging and splitting cells, as non-zero mass fluxes between cells can lead to fluctuations in cell masses. In Arkenstone, this is supplemented with an additional refinement scheme, as described in Section 2.4.
@@
We perform a number of numerical experiments using Arkenstone-Hot in cosmological box simulations to explore the ability of high specific energy winds to act as preventative feedback. This is motivated by recent analytic and semi-analytic works suggesting an important role for such winds (Carr et al., 2023; Pandya et al., 2023; Voit et al., 2024a, b), but until now such a scenario has not been investigated in full cosmological hydrodynamical simulations. To study the impact of these winds, we vary the input parameters of the wind – specifically the mass and energy loadings – and investigate changes in the population-level properties of galaxies due to these variations. We emphasise that we have not calibrated the input parameters of the Arkenstone wind model in this work to match observations, we simply test the impact of certain parameter choices motivated by analytic and semi-analytic modelling, and high resolution ISM simulations.
As a comparison we present our results alongside those from the existing TNG model (which is calibrated to match certain observations), described below, which uses a much more ejective feedback model to successfully reproduce a whole host of galaxy observables. Where available, we also show observationally inferred data for galaxy properties. We note, however, that in this work we do not forward model our simulation predictions into the observational space, so direct comparisons with observational data should be done with caution.
@@
The cosmological simulations in this work all use the same periodic box with a comoving side length of $36.9$ Mpc. The box has a cosmology consistent with Planck’s 2015 results (Planck Collaboration et al., 2016): $\Omega_{\mathrm{m}}=0.3089$, $\Omega_{\Lambda}=0.6911$, $\Omega_{\mathrm{b}}=0.0486$, $h=0.6774$, $\sigma_{8}=0.8159$, $n_{\mathrm{s}}=0.9667$. Initial conditions were generated with the N-GenIC code (Springel et al., 2005; Springel, 2015).
@@
The box contains $512^{3}$ dark matter particles and is initialised with the same number of gas cells, corresponding to a dark matter particle mass of $1.24\times 10^{7}\,\mathrm{M}_{\odot}$ and a base resolution baryon mass of $2.32\times 10^{6}\,\mathrm{M}_{\odot}$. Gravitational softening lengths for DM and stars are 1.48 ckpc at $z>1$, and 0.74 pkpc at $z\leq 1$. Gas softenings are adaptive, but have a minimum of 0.37 pkpc. This is typical of large-scale cosmological boxes like TNG100 (Springel et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018; Pillepich et al., 2018b), EAGLE (Schaye et al., 2015; Crain et al., 2015), Horizon-AGN (Dubois et al., 2014), and Simba (Davé et al., 2019).
@@ 2.2 Galaxy formation model
Aside from the supernova–driven wind model, all simulations presented in this paper share the same physical model in every respect. This model is the same as the fiducial IllustrisTNG simulations, with the exception of magnetohydrodynamics (MHD), that we omit in this work [^3] The Arkenstone model is compatible with MHD, but for simplicity in these first simulations we use hydrodynamics.. We will describe the Arkenstone wind model in Sections 2.3 and 2.4, and the TNG wind model in Section 2.5. First, however, we describe the physics common to all simulations (see Vogelsberger et al., 2013; Weinberger et al., 2017; Pillepich et al., 2018a, for further details).
@@
The resolution achievable in the ISM within cosmological simulations precludes the explicit modelling of a multiphase medium. We therefore adopt a modified version of the effective equation of state (eEoS) model from Springel & Hernquist (2003) to capture the unresolved, local effects of feedback on the ISM. The modifications are the same as in Illustris and TNG (see Vogelsberger et al., 2013); to avoid overpressurising the ISM, we interpolate between the Springel & Hernquist (2003) eEoS and an isothermal EoS at 104 K, with a fraction of 0.3 contributed by the eEoS. All gas cells with a number density above a threshold $n_{\rm H,0}\approx 0.1\,\mathrm{cm}^{-3}$ become part of the eEoS. Metal enrichment occurs via the release of metals to neighbouring cells of star particles in the ISM, at a rate determined by stellar population ages (see Vogelsberger et al., 2013; Pillepich et al., 2018a).
@@
We use the model for BH seeding, growth, and feedback from TNG unchanged (for full details, see Weinberger et al., 2017). BHs of mass $1.18\times 10^{6}\,$ are seeded into haloes with a mass larger than $7.38\times 10^{10}\,$. They can grow via mergers, or through accreting gas at the Eddington-capped Bondi accretion rate (Hoyle & Lyttleton, 1939; Bondi & Hoyle, 1944; Bondi, 1952). Feedback energy is injected as a thermal dump local to the BH at high accretion rates, and as kinetic ‘kicks’ in the low accretion rate regime, motivated by expectations for quasar-driven winds and advection-dominated inflow-outflow solution (ADIOS) winds, respectively. We note that the interaction between Arkenstone and the TNG BH accretion and feedback models will not be trivial, as the parameters of the BH model were calibrated alongside the original TNG wind model. The different stellar feedback prescriptions may therefore affect BH growth differently. For this reason we restrict the bulk of our analysis in this paper to halo masses for which black holes do not strongly affect the evolution of galaxies (below halo masses of $\sim\!10^{12.5}\,$), and we defer a detailed study of the combined feedback effect to future work.
@@
We also use a friends-of-friends algorithm (Davis et al., 1985) and the Subfind halo finder (Springel et al., 2001; Dolag et al., 2009) to identify haloes and galaxies on-the-fly.

### v2

@@ Abstract
We deploy the new Arkenstone galactic wind model in cosmological simulations for the first time, allowing us to robustly resolve the evolution and impact of high specific energy winds. In a (25 $h^{-1}$ Mpc)3 box we perform a set of numerical experiments that systematically vary the mass and energy loadings of such winds, finding that their energy content is the key parameter controlling the stellar to dark matter mass ratio. Increasing the mass loading, at fixed energy, actually results in mildly enhanced star formation, counter to prevailing wisdom, due to the wind becoming cooler. Of the simple parametrisations that we test, we find that an energy loading that scales inversely with halo mass best matches a wide range of observations and can do so with mass loadings drastically lower than those in most previous cosmological simulations. In this scenario, much less material is ejected from the interstellar medium. Instead, winds both heat gas in the circumgalactic medium, slowing infall onto the galaxy, and also drive shocks beyond the virial radius, decreasing the halo-scale accretion rate. We can also report that a much lower fraction of the available supernova energy is needed in preventative galaxy regulation than required by ejective wind feedback models such as IllustrisTNG. This is a Learning the Universe collaboration publication.
@@
pubyear: 2025pagerange: Prevention is better than cure? Feedback from high specific energy winds in cosmological simulations with ArkenstoneA
@@ 1 Introduction
[…]
@@
Feedback is generally believed to come from two primary sources: black holes (BHs) and stars (see e.g. Heckman & Best, 2023). In lower mass haloes, below approximately $10^{12}$ M, the dominant effect is thought to be stellar feedback, from stellar winds, radiation, and supernovae, while massive galaxies are regulated by BH feedback (see e.g. Somerville & Davé, 2015). At all mass scales, regulation of star formation occurs ultimately by reducing the efficiency of the conversion of gas into stars. On small spatial scales, energy injection causes heating and drives turbulence in the ISM, disrupting star formation (e.g., Ostriker et al., 2010; Hopkins et al., 2011; Faucher-Giguère et al., 2013; Krumholz et al., 2018; Ostriker & Kim, 2022). The key to the long-term regulation of star formation however lies in the larger scale baryon cycle, in which galactic winds play a key role (e.g., Strickland & Stevens, 2000; Oppenheimer & Davé, 2008; Hopkins et al., 2012; Anglés-Alcázar et al., 2017; Wright et al., 2024).
@@
Winds can regulate galaxy formation in two ways. They can remove material from the ISM, reducing the amount of gas available to form stars and carrying mass and metals out into the circumgalactic medium (CGM) – ejective feedback. They can also slow or stop gas from accreting onto the galaxy to replenish the fuel for star formation, via the injection of energy into the CGM (or beyond) – preventative feedback. We can characterise winds with two key parameters: the mass loading ($\eta_{\rm M}$, the ratio of the wind mass flux and the star formation rate) and the energy loading ($\eta_{\rm E}$, the ratio of the wind energy flux and the rate at which energy is released during star formation and stellar evolution). We will discuss these parameters extensively in this work.
@@
A wealth of observational evidence exists for outflowing material around many galaxies (see Rupke, 2018, for a review), including from spectroscopic detections of outflows themselves and metal pollution of the CGM at large distances from the central galaxy (e.g. Prochaska et al., 2017; Guo et al., 2020; Zabl et al., 2021). These outflows are ubiquitously found to be multiphase, with molecular, neutral and ionised gas found to co-exist using probes across the electromagnetic spectrum (e.g. Rupke, 2018; Thompson & Heckman, 2024). Despite the widespread detection of outflows around galaxies, estimating the mass and energy loading values is very difficult – especially in the hottest gas phases. Where such data exist, mass fluxes are reliably found to be comparable to or below the star formation rate (e.g. McQuinn et al., 2019; Marasco et al., 2022; Kado-Fong et al., 2024) and the energy flux is found to be comparable to the energy injected by supernovae (e.g. Strickland & Heckman, 2009). Though most observational estimates of mass loading are based on nearby galaxies, there have also been a number of recent studies of outflows at cosmic noon and above, which suggest a similar picture, but with differences across different gas phases and galaxy masses at higher redshift (Concas et al., 2022; Cooper et al., 2025; Xu et al., 2025).
@@
Most subgrid models launch a single wind phase, with the same launch velocity and mass loading (e.g. Vogelsberger et al., 2013; Pillepich et al., 2018a), though several also add a thermal component to either all wind particles (Pillepich et al., 2018a) or a fraction of particles (Davé et al., 2016, 2019) to mimic the impact of hot wind material. No cosmological simulation reliably resolves the interaction between hot and cold phases of material in a wind, which would likely require sub-solar mass resolution (e.g., McCourt et al., 2018; Gronke & Oh, 2020; Abruzzo et al., 2024). An additional class of simulations attempts to resolve more details of the multiphase nature of the ISM in cosmological zoom simulations and to self-consistently track the interaction of stellar feedback and the ISM in launching outflows (e.g. Hopkins et al., 2014, 2018; Muratov et al., 2015; Marinacci et al., 2019; Pandya et al., 2021). Such simulations are difficult to scale up to large cosmological volumes, though some have been used to calibrate loadings in larger-scale simulations (e.g. Davé et al., 2016).
@@
Most modern cosmological box simulations tune the parameters of their models to match particular observables (e.g. the stellar mass – halo mass relation). After this process, the mass and energy loading factors for stellar feedback in most cosmological simulations are typically much higher than what is measured in resolved ISM simulations (e.g. Li & Bryan, 2020). Different simulations can have drastically different predictions for secondary properties of galaxies, most notably in the CGM (Fielding et al., 2017; Davies et al., 2019, 2020; Fielding et al., 2020; Kelly et al., 2022; Ayromlou et al., 2023; Wright et al., 2024). Resolution also affects the properties of the simulated CGM, further complicating the matter, though some recent progress has been made using simulations that specifically focus computational resources by adding additional resolution elements in the CGM (e.g. Hummels et al., 2019; Peeples et al., 2019; van de Voort et al., 2019; Suresh et al., 2019; Bennett & Sijacki, 2020; Ramesh & Nelson, 2024). How many of the differences in predicted galaxy and CGM properties between simulations are due to the numerical implementation of subgrid models is unclear; until now it has not been possible to study the regulatory effect of high specific energy, low mass winds on a large population of galaxies.
@@
Simulating high specific energy winds is challenging in cosmological simulations (see section 2 of Smith et al., 2024a, for a detailed discussion). Hot, fast winds with low mass fluxes have lower densities, leading to coarser spatial resolution in a Lagrangian code (or with a quasi-Lagrangian refinement strategy in an Eulerian scheme, as is typical in cosmological volumes). Even with the highest mass resolution of current cosmological box simulations ($\sim\!10^{5}$ M) such a wind can become completely unresolved, with single resolution elements being larger than the galaxy launching them (see fig. 1 of Smith et al., 2024a).
@@
Arkenstone, a new subgrid model for multiphase winds, was developed to address these issues (Smith et al., 2024a, b). The full model launches winds containing both a high specific energy component (Arkenstone-Hot, Smith et al., 2024a, henceforth S24) and an unresolved cold cloud component (Arkenstone-Cold, Smith et al., 2024b), each with their own mass and energy loadings. Arkenstone is designed to capture the effects of multiphase winds in large-scale cosmological simulations, with resolutions coarser than some of the high-resolution zooms that attempt to model such outflows self-consistently. We note that even for such zooms it is currently impossible to capture the interaction between wind phases, a key aspect of the full Arkenstone model. With this full model, we can explore a huge area of previously unexplored parameter space of mass and energy loadings to investigate the balance between ejective and preventative feedback in galaxy formation. The ultimate aim is to use parameters calibrated with highly resolved ISM simulations (e.g. Kim et al., 2020b) as inputs to the Arkenstone model, to remove the need to calibrate to global galaxy observations, as is commonly done currently in most cosmological simulations. This is central to the goals of the SMAUG [^1] https://www.simonsfoundation.org/flatiron/center-for-computational-astrophysics/galaxy-formation/smaug/ (Simulating Multiscale Astrophysics to Understand Galaxies) and Learning the Universe [^2] http://learning-the-universe.org collaborations.
@@
In this first paper applying the Arkenstone model to cosmological simulations we focus only on the hot component of winds, allowing us to study galaxy formation in a regime where preventative feedback dominates and little mass is ejected from the ISM. We vary the key parameters of these hot winds – mass and energy loadings – to study how the properties of the galaxy population respond. The paper is structured as follows: In Section 2 we describe the Arkenstone model and the simulations in this paper. Sections 3.1 and 3.2 show properties of the galaxies and their gaseous haloes in our simulations at $z=0$. Section 3.4 then looks at how the feedback from Arkenstone regulates star formation through preventative, rather than ejective feedback. We look at the redshift evolution of star formation in Section 3.5, discuss comparisons to recent analytic, semi-analytic and simulation work in Section 4, and summarise our results in Section 5.
@@ 2.1 Simulation set-up
The simulations in this paper use Arepo (Springel, 2010), a cosmological hydrodynamical code that incorporates gravity using a TreePM scheme and solves the equations of hydrodynamics on a quasi-Lagrangian, unstructured, moving Voronoi mesh, where mesh generating points move with the fluid velocity (with small corrections in selected cells to promote a regular mesh geometry). In Arepo, a default refinement scheme keeps gas cells within a factor of two of a given mass resolution by merging and splitting cells, as non-zero mass fluxes between cells can lead to fluctuations in cell masses. In Arkenstone, this is supplemented with an additional refinement scheme, as described in Section 2.5.
@@
We perform a number of numerical experiments using Arkenstone-Hot in cosmological box simulations to explore the ability of high specific energy winds to act as preventative feedback. To study the impact of these winds, we vary the input parameters of the wind – specifically the mass and energy loadings – and investigate changes in the population-level properties of galaxies due to these variations. We emphasise that we have not calibrated the input parameters of the Arkenstone wind model in this work to match observations, we simply test the impact of certain parameter choices motivated by analytic and semi-analytic modelling, and high resolution ISM simulations.
@@
The cosmological simulations in this work all use the same initial conditions - a periodic box with a comoving side length of $36.9$ Mpc. The box has a cosmology consistent with the Planck 2015 constraints (Planck Collaboration et al., 2016): $\Omega_{\mathrm{m}}=0.3089$, $\Omega_{\Lambda}=0.6911$, $\Omega_{\mathrm{b}}=0.0486$, $h=0.6774$, $\sigma_{8}=0.8159$, $n_{\mathrm{s}}=0.9667$. Initial conditions were generated with the N-GenIC code (Springel et al., 2005; Springel, 2015).
@@
The simulations contain $512^{3}$ dark matter particles and are initialised with the same number of gas cells, corresponding to a dark matter particle mass of $1.24\times 10^{7}\,\mathrm{M}_{\odot}$ and a base resolution baryon mass of $2.32\times 10^{6}\,\mathrm{M}_{\odot}$. As described in Section 2.4, wind particles and flagged cells in Arkenstone have a resolution 100 times better than the base resolution, corresponding to $2.32\times 10^{4}\,\mathrm{M}_{\odot}$. Gravitational softening lengths for DM and stars are 1.48 ckpc at $z>1$, and 0.74 pkpc at $z\leq 1$. Gas softenings are adaptive, but have a minimum of 0.37 pkpc. This is typical of large-scale cosmological boxes like TNG100 (Springel et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018; Pillepich et al., 2018b), EAGLE (Schaye et al., 2015; Crain et al., 2015), Horizon-AGN (Dubois et al., 2014), and Simba (Davé et al., 2019).
@@ 2.2 Galaxy formation model
Aside from the supernova–driven wind model, all simulations presented in this paper share the same physical model in every respect. This model is the same as the fiducial IllustrisTNG simulations, with the exception of magnetohydrodynamics (MHD), which is omitted in this work [^3] The Arkenstone model is compatible with MHD, but for simplicity in these first simulations we use hydrodynamics only.. We will describe the Arkenstone wind model in Sections 2.4 and 2.5, and the TNG wind model in Section 2.3. First, however, we describe the physics common to all simulations (see Vogelsberger et al., 2013; Weinberger et al., 2017; Pillepich et al., 2018a, for further details).
@@
The resolution achievable in the ISM within cosmological simulations precludes the explicit modelling of a multiphase medium. We therefore adopt a modified version of the effective equation of state (eEoS) model from Springel & Hernquist (2003) to capture the unresolved, local effects of feedback on the ISM. The modifications are the same as in Illustris and TNG (see Vogelsberger et al., 2013); to avoid overpressurising the ISM, we interpolate between the Springel & Hernquist (2003) eEoS and an isothermal EoS at 104 K, with a thirty per cent contribution from the eEoS. All gas cells with a number density above a threshold $n_{\rm H,0}\approx 0.1\,\mathrm{cm}^{-3}$ join the eEoS. We note that while the pressurisation of the ISM uses a portion of the energy nominally available from supernovae, the uncertainties in both the Springel & Hernquist (2003) model and total amount of energy that supernovae can actually provide mean that, as in TNG, we treat this separately from the energy used to launch winds. As we discuss in Section 2.6, Arkenstone uses considerably less energy than TNG to launch winds, and uses a total energy well within plausible limits for the energy released from supernovae. Metal enrichment occurs via the release of metals to neighbouring cells of star particles in the ISM, at a rate determined by stellar population ages (see Vogelsberger et al., 2013; Pillepich et al., 2018a, and the references therein).
@@
We use the same model for BH seeding, growth, and feedback as in TNG without modification (for full details, see Weinberger et al., 2017). BHs of mass $1.18\times 10^{6}\,$M are seeded into haloes with a mass larger than $7.38\times 10^{10}\,$M. They can grow via mergers, or through accreting gas at the Eddington-capped Bondi accretion rate (Hoyle & Lyttleton, 1939; Bondi & Hoyle, 1944; Bondi, 1952). Feedback energy is injected as a thermal dump local to the BH at high accretion rates, and as kinetic ‘kicks’ in the low accretion rate regime, motivated by expectations for quasar-driven winds and advection-dominated inflow-outflow solution (ADIOS) winds, respectively. We note that the interaction between Arkenstone and the TNG BH accretion and feedback models will not be trivial, as the parameters of the BH model were calibrated alongside the original TNG wind model. The different stellar feedback prescriptions may therefore affect BH growth. For this reason we restrict the bulk of our analysis in this paper to halo masses for which black holes do not strongly affect the evolution of galaxies (below halo masses of $\sim\!10^{12.5}\,$M), and we defer a detailed study of the combined feedback effect to future work.
@@
As in TNG, we use a friends-of-friends algorithm (Davis et al., 1985) and the Subfind halo finder (Springel et al., 2001; Dolag et al., 2009) to identify haloes and galaxies on-the-fly.
2.3 TNG winds
We now summarise the key parts of the TNG wind model (fully described in Pillepich et al., 2018a). This model uses hydrodynamically decoupled wind particles, launched isotropically from star forming gas cells. TNG uses a density threshold for recoupling (5% of the star formation threshold), at which point the wind particle dumps all of its mass, energy, momentum and metals into its host cell (referred to as simple recoupling in the Arkenstone scheme).
Unlike in Arkenstone, the wind particles in TNG radiatively cool while hydro-decoupled, with their effective density adopted from the nearest gas cell. We omit this in Arkenstone for three reasons. Firstly, for high specific energy winds (which Arkenstone-Hot is designed to treat), the density of the local ISM, as measurable in the simulation, is not relevant – we are implicitly assuming that the hot, fast wind is propagating through unresolved low density channels and superbubble breakouts. Secondly, for these hot winds, the radiative cooling time is too long for this additional cooling to be important. Finally, our philosophy with Arkenstone is to avoid modelling the evolution of the wind within the ISM as much as possible, instead focusing on the properties of the wind outside of the ISM; radiative cooling while in the ISM could be folded into the input energy loadings if desired.
The mass loading is then set as,
$\eta_{\rm M}=\frac{2}{(\Delta v_{\rm w,TNG})^{2}}\left(1-\tau_{\rm w}\right)\eta_{\rm E}u_{\star},$ (2)
where $\tau_{\rm w}=0.1$ is the thermal fraction of the wind energy. Note that here we have expressed Equation 2 with a similar parametrisation to our Equations 4 and 5 in the subsequent section. The combination $u_{\star}\eta_{\rm E}$ is equivalent to the quantity $e_{w}$ in equation 3 of Pillepich et al. (2018a), which has a metallicity dependence (higher wind energy at lower metallicity). Pillepich et al. (2018a) also assumes a SN energy per unit stellar mass formed that is 13% higher than our reference $u_{\star}$ (relating to minor differences in IMF cutoff masses). Since $u_{\star}$ and $\eta_{\rm E}$ are degenerate as far as the actual wind energy input is concerned (and the difference is slight), for the purpose of comparing energy loadings between Arkenstone and TNG on an even footing, we fix $u_{\star}$ here to our value and absorb the difference into the energy loading (along with the metallicity dependence).
The TNG model initially forms wind particles with the same metallicity as their birth cell, but then removes metal mass from wind particles and redistributes this among neighbouring ISM cells before being kicked, preventing the removal of too many metals due to TNG’s high mass loading (see Section 2.6 and Appendix A).